Secret Message - 2

Given a string, write a program to print a secret message that replaces characters with numbers 'a' with 1, 'b' with 2, ..., 'z' with 26 where characters are separated by '-'.


Note: You need to replace both uppercase and lowercase characters. You can ignore replacing all characters that are not letters.


a b c d e f g h i j k l m n o p q r s t u v w x y z

1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26


Input

Explanation


For example, if the given input is "python", "p" should replaced with "16", similarly"y" with "25","t" with "20","h" with "8","o" with "15","n" with "14". So the output should be "16-25-20-8-15-14".

Sample Input 1

python


Sample Output 1

16-25-20-8-15-14


Sample Input 2

Foundations


Sample Output 2

6-15-21-14-4-1-20-9-15-14-19




1
Expert's answer
2021-03-21T23:30:52-0400
def manipulateString(str):

    for i in range(len(str)):
        asc = ord(str[i])

        rem = asc - ( 26-(ord(str[i]) - ord('a')))
        m = rem % 26
        str[i] = m+ord('a')
    
    print(str)
    #print(''.join(str))



# Driver code
if __name__ == "__main__":
    str = "Hello Student"
    str = list(str)
    manipulateString(str)